# Entropic (Gromov) Wasserstein Flow Matching with GENOT

### v3

@@ Abstract
Optimal transport (OT) theory has reshaped the field of generative modeling: Combined with neural networks, recent Neural OT (N-OT) solvers use OT as an inductive bias, to focus on “thrifty” mappings that minimize average displacement costs. This core principle has fueled the successful application of N-OT solvers to high-stakes scientific challenges, notably single-cell genomics. N-OT solvers are, however, increasingly confronted with practical challenges: while most N-OT solvers can handle squared-Euclidean costs, they must be repurposed to handle more general costs; their reliance on deterministic Monge maps as well as mass conservation constraints can easily go awry in the presence of outliers; mapping points across heterogeneous spaces is out of their reach. While each of these challenges has been explored independently, we propose a new framework that can handle, natively, all of these needs. The generative entropic neural OT (GENOT) framework models the conditional distribution $\pi_{\varepsilon}(\mathbf{y}|\mathbf{x})$ of an optimal entropic coupling $\pi_{\varepsilon}$, using conditional flow matching. GENOT is generative, and can transport points across spaces, guided by sample-based, unbalanced solutions to the Gromov-Wasserstein problem, that can use any cost. We showcase our approach on both synthetic and single-cell datasets, using GENOT to model cell development, predict cellular responses, and translate between data modalities.
Machine Learning, Optimal Transport
@@ 1 Introduction
Mapping a probability distribution onto another is a ubiquitous challenge in machine learning, with many implications in generative modeling. Optimal transport (OT) has arisen in a few years as a major purveyor of tools to better address these challenges, both in theory and practice. OT focuses on finding maps that can effectively transform a distribution of matter onto another, by minimizing a certain notion of cost (Santambrogio, 2015). Applying OT to large-dimensional problems arising in machine learning and sciences has necessitated various modifications and adaptations. Starting with solvers that can solve approximate matching problems at large scales (Cuturi, 2013; Peyré et al., 2016; Scetbon et al., 2021, 2022), a recent plethora of OT-inspired training approaches for neural networks have been proposed (Makkuva et al., 2020; Korotin et al., 2020; Asadulaev et al., 2022; Fan et al., 2020; Uscidda & Cuturi, 2023; Lipman et al., 2023; Tong et al., 2020, 2023b). Mirroring these new possibilities in the field of single-cell genomics, arguably one of the most exciting transfers of OT to science, the use of OT has evolved from point-wise matching problems (Schiebinger et al., 2019; Demetci et al., 2022), to neural-based approaches that can, for instance, predict the response of cells to various perturbations (Bunne et al., 2021, 2022). As a testament of their success, practitioners demand ever more flexibility from Neural-OT solvers, in a suite of increasingly complex challenges:
From Deterministic to Stochastic Maps. The classic OT pipeline rests on deterministic (Monge) maps. Such maps lack flexibility, both at estimation and inference time. Practitioners may favor, instead, stochasticity, which would account naturally for instance, for the non-determinism of cell evolutions (Elowitz et al., 2002). Stochastic formulations can also produce a conditional distribution that can be used to quantify uncertainty. In the discrete setting, this property is fulfilled by entropy-regularized OT (EOT) (Cuturi, 2013). In a neural setting, some works (Korotin et al., 2022, 2023b; Asadulaev et al., 2022) have explored computing weak OT coupling by solving a min-max problem.
From Linear to Quadratic Problems. Classic “linear” OT puts in correspondence distributions defined in the same space $\mathcal{X}$, using a cost $c(\mathbf{x},\mathbf{y})$ for any two points $\mathbf{x},\mathbf{y}\in\mathcal{X}$ (most often set to be the squared Euclidean distance (Santambrogio, 2015, §1.3) in applications). Yet, more challenging use-cases arising in practice now involve source and target distributions that do not live in the same space, e.g. $\mathcal{X}$ and $\mathcal{Y}$ have differing dimensions, as in Demetci et al. (2022). The challenge, in that case, lies in the absence of a “natural” cost function between vectors of different sizes. This can be bypassed using quadratic losses (Mémoli, 2011; Sturm, 2020), a.k.a. Gromov-Wasserstein (GW) problems. Despite the challenges inherent to solving GW, practitioners have expressed much interest in that flexibility, as evidenced by the wide adoption of the mixed linear/quadratic fused Gromov-Wasserstein (FGW) distance (Vayer et al., 2018), see e.g. Klein et al. (2023); Lange et al. (2023); Nitzan et al. (2019); Zeira et al. (2022). To our knowledge, the only neural formulation for GW proposed thus far by (Nekrashevich et al., 2023) learns deterministic maps for the inner product costs, using a min-max-min optimization procedure.
Flexibility in Mass Conservation. In many real-world applications, the data acquisition process can be error-prone, resulting in outliers. To mitigate this, unbalanced OT (UOT) formulations that can discard observations have been proposed (Frogner et al., 2015; Chizat et al., 2018; Séjourné et al., 2021), with numerous applications to generative modeling (Balaji et al., 2020; Yang & Uhler, 2019; Eyring et al., 2023) and single-cell genomics (Schiebinger et al., 2019; Eyring et al., 2022; Lübeck et al., 2022).
@@
Contributions. We propose a flexible neural OT framework that satisfies all the requirements above:
@@
  • We propose GENOT, the first method that parameterizes linear and quadratic EOT couplings by modeling their conditional distributions with neural networks (Prop. 3.1), using flow matching (Lipman et al., 2023) as a core block, and any cost function between samples.
  • By showing that an unbalanced EOT problem is equivalent to a balanced one between re-weighted measures (Prop. 3.2), that can be estimated consistently (Prop. 3.3), we introduce U-GENOT to solve unbalanced problems.
@@
  • We extend (U-)GENOT to solve the fused problem (§ 3.3). To our knowledge, GENOT is the first method to solve a continuous fused problem.
  • We show that GENOT outperforms other linear neural EOT solvers on the Gushchin et al. (2023) benchmark,
@@
  • We showcase GENOT’s ability to handle crucial single-cell biology challenges: we (i) quantify lineage branching events in the developing mouse pancreas, (ii) predict cellular responses to drug perturbations, along with a well-calibrated uncertainty estimation, and (iii) introduce a novel method to translate ATAC-seq to RNA-seq data.
@@ 2 Background
Notations. We consider in this work two compact sets $\mathcal{X}\subset\mathbb{R}^{p}$, $\mathcal{Y}\subset\mathbb{R}^{q}$, referred to as the source and the target domain, respectively. In general,  $p\neq q$. The set of positive (resp. probability) measures on $\mathcal{X}$ is denoted by $\mathcal{M}^{+}(\mathcal{X})$ (resp. $\mathcal{M}_{1}^{+}(\mathcal{X})$). For $\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})$, $\pi_{1}:=p_{1}\sharp\pi$ and $\pi_{2}:=p_{2}\sharp\pi$ denote its marginals. Then, for $\mu\in\mathcal{M}^{+}(\mathcal{X}),\nu\in\mathcal{M}^{+}(\mathcal{Y})$,  $\Pi(\mu,\nu)=\{\pi:\,\pi_{1}=\mu,\,\pi_{2}=\nu\}$. Moreover, $\tfrac{\mathop{}\!\mathrm{d}\mu}{\mathop{}\!\mathrm{d}\nu}$ denotes the relative density of $\mu$ w.r.t. $\nu$, and we write $\mu=\tfrac{\mathop{}\!\mathrm{d}\mu}{\mathop{}\!\mathrm{d}\nu}\cdot\nu$. For $\rho,\gamma\in\mathcal{M}^{+}(\mathcal{X})$, $\mathrm{KL}(\rho|\gamma)=\int_{\mathcal{X}}\log(\frac{\mathop{}\!\mathrm{d}\rho}{\mathop{}\!\mathrm{d}\gamma})\mathop{}\!\mathrm{d}\rho-\int_{\mathcal{X}}\mathop{}\!\mathrm{d}\gamma+\int_{\mathcal{X}}\mathop{}\!\mathrm{d}\rho$.
Linear Entropic OT. Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ be a cost function, $\mu\in\mathcal{M}_{1}^{+}(\mathcal{X}),\nu\in\mathcal{M}_{1}^{+}(\mathcal{Y})$ and $\varepsilon\geq 0$. The linear entropy-regularized OT problem reads
@@
(LEOT) is also known as the static Schrödinger bridge (SB) problem (Léonard, 2013). With $\varepsilon=0$, we recover the Kantorovich (1942) problem. When $\mu$ and $\nu$ are discrete, we can solve (LEOT) with the Sinkhorn algorithm (Cuturi, 2013). For measures on $n$ points, its complexity is $\mathcal{O}(n^{2})$ in time, and $\mathcal{O}(n)$ or $\mathcal{O}(n^{2})$ in memory, depending on $c$.
@@
Quadratic Entropic OT. As opposed to considering an inter-domain cost defined on $\mathcal{X}\times\mathcal{Y}$, quadratic entropic OT is concerned with seeking couplings based on intra-domain cost functions $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$:
@@
$\min_{\pi\in\Pi(\mu,\nu)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}$ $Q_{c_{\mathcal{X}},c_{\mathcal{Y}}}\mathop{}\!\mathrm{d}(\pi\otimes\pi)+\varepsilon\,\text{KL}(\pi\|\mu\otimes\nu),$ (QEOT)
@@
where $Q_{c_{\mathcal{X}},c_{\mathcal{Y}}}(\mathbf{x},\mathbf{y},\mathbf{x}^{\prime},\mathbf{y}^{\prime}):=|c_{X}(\mathbf{x},\mathbf{x}^{\prime})-c_{Y}(\mathbf{y},\mathbf{y^{\prime}})|^{2}$ measures the pointwise cost distortion. With $\varepsilon=0$, we recover the Gromov-Wasserstein (Mémoli, 2011) problem. This is the standard OT formulation for measures supported on incomparable spaces. Using $\varepsilon>0$ offers computational advantages, since for discrete $\mu$, $\nu$, we can solve  (QEOT) with a mirror-descent scheme iterating the Sinkhorn algorithm (Peyré et al., 2016). This solver has $\mathcal{O}(n^{2})$ or $\mathcal{O}(n^{3})$ time complexity, depending on $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (Scetbon et al., 2022, Alg. 1 & Alg. 2), and $\mathcal{O}(n^{2})$ memory complexity.
@@
Unbalanced Extensions. The EOT formulations presented above can only handle measures with the same total mass. Unbalanced optimal transport (UOT) (Liero et al., 2018; Chizat et al., 2018) lifts this constraint by penalizing the deviation of $p_{1}\sharp\pi$ to $\mu$ and $p_{2}\sharp\pi$ to $\nu$ with a divergence. Using the $\mathrm{KL}$ and introducing weightings $\lambda_{1},\lambda_{2}>0$ the unbalanced extension of (LEOT) reads
@@
$\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}$ $\int_{\mathcal{X}\times\mathcal{Y}}c\mathop{}\!\mathrm{d}\pi+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)$ (ULEOT)
$+\lambda_{1}\mathrm{KL}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}(\pi_{2}|\nu).$
@@
This problem can be solved efficiently in the discrete setting using a variant of the Sinkhorn algorithm (Frogner et al., 2015; Séjourné et al., 2023a). Analogously, quadratic OT also admits an unbalanced generalization, which reads
@@
$\begin{split}\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}&\int_{(\mathcal{X}\times\mathcal{Y})^{2}}Q_{c_{\mathcal{X}},c_{\mathcal{Y}}}\mathop{}\!\mathrm{d}(\pi\otimes\pi)\,+\varepsilon\mathrm{KL}^{\otimes}(\pi|\mu\otimes\nu)\\ &+\lambda_{1}\mathrm{KL}^{\otimes}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}^{\otimes}(\pi_{2}|\nu),\end{split}$ (UQEOT)
@@
where $\mathrm{KL}^{\otimes}(\rho|\gamma)=\mathrm{KL}(\rho\otimes\rho|\gamma\otimes\gamma)$. This can be solved using an extension of Peyré et al. (2016)’s scheme introduced by Séjourné et al. (2023b). Each solver has the same time and memory complexity as its balanced counterpart. For both (ULEOT) and (UQEOT), instead of selecting $\lambda_{i}$, we introduce $\tau_{i}=\tfrac{\lambda_{i}}{\lambda_{i}+\varepsilon}$ s.t. we recover the hard marginal constraint for $\tau_{i}=1$, when $\lambda_{i}\rightarrow+\infty$. We write $\tau=(\tau_{1},\tau_{2})$.
@@
Flow Matching. Given a prior $\rho_{0}\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$ and a time-dependent vector field $(v_{t})_{t\in[0,1]}$, one can define a probability path $p_{t}$ starting from $\rho_{0}$ using the flow $\phi_{t}$ solving
@@
by setting $p_{t}=\phi_{t}\sharp\rho_{0}$. We then say that $v_{t}$ generates the path $p_{t}$. Continuous Normalizing Flows (Chen et al., 2018) (CNFs) model $v_{t,\theta}$ with a neural network , which is trained to match a terminal condition $p_{1}=\rho_{1}\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$. Conditional Flow Matching (CFM) (Lipman et al., 2023) is a simulation-free technique to train CNFs by constructing individual paths between samples, and minimizing
@@
where $[Z_{0},Z_{1}]_{t}:=(1-t)Z_{0}+tZ_{1}$ with $t\sim\mathcal{U}([0,1])$. If this loss is 0, the flow maps $\rho_{0}$ to $\rho_{1}$, i.e., $\phi_{1}\sharp\rho_{0}=\rho_{1}$.

### v4

# GENOT: Entropic (Gromov) Wasserstein Flow Matching with Applications to Single-Cell Genomics

@@ Abstract
Single-cell genomics has significantly advanced our understanding of cellular behavior, catalyzing innovations in treatments and precision medicine. However, single-cell sequencing technologies are inherently destructive and can only measure a limited array of data modalities simultaneously. This limitation underscores the need for new methods capable of realigning cells. Optimal transport (OT) has emerged as a potent solution, but traditional discrete solvers are hampered by scalability, privacy, and out-of-sample estimation issues. These challenges have spurred the development of neural network-based solvers, known as neural OT solvers, that parameterize OT maps. Yet, these models often lack the flexibility needed for broader life science applications. To address these deficiencies, our approach learns stochastic maps (i.e. transport plans), allows for any cost function, relaxes mass conservation constraints and integrates quadratic solvers to tackle the complex challenges posed by the (Fused) Gromov-Wasserstein problem. Utilizing flow matching as a backbone, our method offers a flexible and effective framework. We demonstrate its versatility and robustness through applications in cell development studies, cellular drug response modeling, and cross-modality cell translation, illustrating significant potential for enhancing therapeutic strategies.
@@ 1 Introduction
Discrete Optimal Transport in Single-Cell Genomics. Due to the destructive nature of single-cell sequencing technologies, it is fundamental to realign distributions of sequenced cells. Discrete optimal transport (OT) has proven useful in this task. For example, it is standard practice to leverage the framework of OT to reconstruct cellular trajectories across developmental time points [72] and to study the response of cells to external perturbations like drugs or gene knockouts [18]. In these settings, both distributions, i.e. earlier and later time point and before perturbation and after perturbation, respectively, live in the same space, a setting which we refer to as linear OT. In contrast, most applications of OT in single-cell genomics require maps across (partially) incomparable spaces, the setting which we refer to as quadratic OT. For example, analysis of spatial transcriptomics datasets requires the quadratic formulation of OT as the coordinate systems of two different slices are warped and rotated, thus not living in the same coordinate space [60]. Similarly, adding spatial information as a prior has been shown to improve the performance of modeling the trajectory of cells in the discrete setting [39]. Analogously, adding information from lineage tracing readouts [89] recovers the evolution of cells more faithfully [42]. Another prevalent task in single-cell genomics which is commonly approached with quadratic OT solvers is the alignment of cells across modalities [17]. Most single-cell technologies capture only one modality, which is insufficient to obtain a comprehensive representation of the cellular state. As single-cell datasets grow larger [65], traditional OT solvers become less applicable to tasks in single-cell genomics due to their high computational complexity. Thanks to recent advancements in low-rank OT [69, 70, 71], the aforementioned OT-based algorithms are more accessible to single-cell biologists [39]. Yet, the non-parametric nature of discrete OT solvers entails issues with respect to data privacy and prevents their application to large scale single-cell atlases capturing millions of cells [65]. Moreover, out-of-samples estimation is limited to very specific scenarios [63].
Current Limitations of Neural Optimal Transport in Single-Cell Genomics. To overcome these limitations, neural OT solvers have been leveraged to study cellular perturbations [7, 8, 86] and to model cellular trajectories [83, 21]. Yet, these methods estimate Monge maps, i.e., deterministic maps, contradicting the assumption that cells evolve stochastically  [19]. Stochastic formulations are also favorable as they can produce a conditional distribution that can be used to quantify uncertainty [25]. In the discrete setting, stochastic maps can be obtained from entropy-regularized OT (EOT) [13]. Recently, a number of works have addressed learning EOT plans [85, 84, 29, 76, 41] in the neural setting. Yet, all of these methods are limited in the choice of the cost function, with most of them being restricted to the squared Euclidean cost. Single-cell genomics data is known to be non-Euclidean [57] and thus requires a flexible choice of the cost function in OT applications as demonstrated in the discrete OT setting in Demetci et al. [17], Huguet et al. [35], Klein et al. [39]. The third requirement for applying OT to single-cell genomics is the option to lift the mass conservation constraint, allowing for unbalanced optimal transport [24, 11, 74]. Unbalanced OT is crucial to model cellular growth and death as well as for automatically discarding outliers, which are prevalent in highly noisy measurements in single-cell genomics. Yang and Uhler [94], Lübeck et al. [50], Eyring et al. [20] proposed ways to incorporate unbalancedness into deterministic linear OT maps, while unbalanced formulations for entropic neural OT have barely been explored. The most severe shortcoming of the existing plethora of neural OT estimators is their limitation to the linear OT scenario, i.e., to learning maps within spaces. Yet, most applications in single-cell genomics require the Gromov-Wasserstein (GW) [54] or the fused Gromov-Wasserstein (FGW) [88] formulation. To the best of our knowledge, the only neural formulation for GW proposed thus far learns deterministic, balanced maps for the inner product costs, using a min-max-min optimization procedure, severely limiting its applications [58]. Hence, we arrive at four necessities that need to be fulfilled and have to be flexibly combinable to make neural OT generally applicable to problems in single-cell genomics:
  • N1: modeling the evolution of cells stochastically as opposed to deterministically,
  • N2: flexibly choosing cost functions due to non-Euclidean geometry of single-cell genomics data,
  • N3: allowing for unbalanced OT to model cellular growth and death or accounting for outliers,
  • N4: mapping across completely or partially incomparable spaces.
@@
Contributions. Hence, we propose GENOT (Generative Entropic Neural Optimal Transport), a powerful and flexible neural OT framework that satisfies all of the above requirements:
@@
  • GENOT is the first method that parameterizes linear and quadratic EOT couplings for any cost by modeling their conditional distributions, using flow matching [45] as a backbone.
  • We extend GENOT to the unbalanced setting, resulting in U-GENOT, to flexibly allow for mass variations in both the linear and the quadratic setting.
@@
  • We extend (U-)GENOT to address the fused problem and this way propose, to the best of our knowledge, the first neural OT solver for the FGW problem.
@@
  • We showcase GENOT’s ability to handle common challenges in single-cell biology: we (i) quantify lineage branching events in the developing mouse pancreas, (ii) predict cellular responses to drug perturbations, along with a well-calibrated uncertainty estimation, and (iii) translate ATAC-seq to RNA-seq with GW and introduce a novel method to perform this translation task with FGW.
@@ 2 Background
Notations. Let $\mathcal{X}\subset\mathbb{R}^{p}$, $\mathcal{Y}\subset\mathbb{R}^{q}$ compact sets, referred to as the source and the target domain, respectively. In general,  $p\neq q$. The set of positive (resp. probability) measures on $\mathcal{X}$ is denoted by $\mathcal{M}^{+}(\mathcal{X})$ (resp. $\mathcal{M}_{1}^{+}(\mathcal{X})$). For $\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})$, $\pi_{1}:=p_{1}\sharp\pi$ and $\pi_{2}:=p_{2}\sharp\pi$ denote its marginals. Then, for $\mu\in\mathcal{M}^{+}(\mathcal{X}),\nu\in\mathcal{M}^{+}(\mathcal{Y})$,  $\Pi(\mu,\nu)=\{\pi:\,\pi_{1}=\mu,\,\pi_{2}=\nu\}$. $\tfrac{\mathop{}\!\mathrm{d}\mu}{\mathop{}\!\mathrm{d}\nu}$ denotes the relative density of $\mu$ w.r.t. $\nu$, s.t. $\mu=\tfrac{\mathop{}\!\mathrm{d}\mu}{\mathop{}\!\mathrm{d}\nu}\cdot\nu$. For $\rho,\gamma\in\mathcal{M}^{+}(\mathcal{X})$, $\mathrm{KL}(\rho|\gamma)=\int_{\mathcal{X}}\log(\frac{\mathop{}\!\mathrm{d}\rho}{\mathop{}\!\mathrm{d}\gamma})\mathop{}\!\mathrm{d}\rho-\int_{\mathcal{X}}\mathop{}\!\mathrm{d}\gamma+\int_{\mathcal{X}}\mathop{}\!\mathrm{d}\rho$.
Linear Entropic OT. Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ be a continuous cost function, $\mu\in\mathcal{M}_{1}^{+}(\mathcal{X}),\nu\in\mathcal{M}_{1}^{+}(\mathcal{Y})$ and $\varepsilon\geq 0$. The linear entropy-regularized OT problem reads
@@
A solution $\pi^{\star}_{\varepsilon}$ of (LEOT) always exists and is unique when $\varepsilon>0$. (LEOT) is also known as the static Schrödinger bridge (SB) problem [52]. With $\varepsilon=0$, we recover the Kantorovich [38] problem. For discrete $\mu$ and $\nu$, we can solve (LEOT) with the Sinkhorn algorithm [13], whose complexity for $n$ points is $\mathcal{O}(n^{2})$ in time, and $\mathcal{O}(n)$ or $\mathcal{O}(n^{2})$ in memory, depending on $c$.
@@
Quadratic Entropic OT. As opposed to considering an inter-domain cost defined on $\mathcal{X}\times\mathcal{Y}$, quadratic entropic OT is concerned with seeking couplings that minimize the distortion of the geometries induced by intra-domain cost functions $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$:
@@
$\min_{\pi\in\Pi(\mu,\nu)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}$ $D_{c_{\mathcal{X}},c_{\mathcal{Y}}}\mathop{}\!\mathrm{d}(\pi\otimes\pi)+\varepsilon\,\text{KL}(\pi\|\mu\otimes\nu),$ (QEOT)
@@
where $D_{c_{\mathcal{X}},c_{\mathcal{Y}}}(\mathbf{x},\mathbf{y},\mathbf{x}^{\prime},\mathbf{y}^{\prime}):=|c_{X}(\mathbf{x},\mathbf{x}^{\prime})-c_{Y}(\mathbf{y},\mathbf{y^{\prime}})|^{2}$ quantifies the pointwise cost distortion. A solution $\pi^{\star}_{\varepsilon}$ to Prob.(QEOT) always exists (see B.1). With $\varepsilon=0$, we recover the Gromov-Wasserstein [54] problem, which is the standard OT formulation for transporting measures supported on incomparable spaces. In addition to statistical benefits [95], using $\varepsilon>0$ also offers computational benefits, since for discrete $\mu$ and $\nu$, we can solve  (QEOT) with a mirror descent scheme iterating the Sinkhorn algorithm [62]. For measures on $n$ points, its time complexity is $\mathcal{O}(n^{2})$ or $\mathcal{O}(n^{3})$, depending on $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ [70, Alg. 1&2], while its memory complexity is always $\mathcal{O}(n^{2})$.
@@
Unbalanced Extensions. The EOT formulations presented above can only handle measures with the same total mass. Unbalanced optimal transport (UOT) [44, 11] lifts this constraint by penalizing the deviation of $p_{1}\sharp\pi$ to $\mu$ and $p_{2}\sharp\pi$ to $\nu$ with a divergence. Using the $\mathrm{KL}$ and introducing weightings $\lambda_{1},\lambda_{2}>0$ the unbalanced extension of (LEOT) reads
@@
$\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}\int_{\mathcal{X}\times\mathcal{Y}}c\mathop{}\!\mathrm{d}\pi+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)+\lambda_{1}\mathrm{KL}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}(\pi_{2}|\nu).$ (ULEOT)
@@
This problem can be solved efficiently in the discrete setting using a variant of the Sinkhorn algorithm [24, 80]. Analogously, quadratic OT also admits an unbalanced generalization, which reads
@@
$\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}D_{c_{\mathcal{X}},c_{\mathcal{Y}}}\mathop{}\!\mathrm{d}(\pi\otimes\pi)\,+\varepsilon\mathrm{KL}^{\otimes}(\pi|\mu\otimes\nu)+\lambda_{1}\mathrm{KL}^{\otimes}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}^{\otimes}(\pi_{2}|\nu),$ (UQEOT)
@@
where $\mathrm{KL}^{\otimes}(\rho|\gamma)=\mathrm{KL}(\rho\otimes\rho|\gamma\otimes\gamma)$. A solution $\pi^{\star}_{\varepsilon,\tau}$ to Prob.(UQEOT) always exists (see B.1). In the discrete setting, Prob. (UQEOT) can be solved using an extension of Peyré et al. [62]’s scheme introduced by Séjourné et al. [81]. Each solver has the same time and memory complexity as its balanced counterpart. For both (ULEOT) and (UQEOT), instead of selecting $\lambda_{i}$, we introduce $\tau_{i}=\tfrac{\lambda_{i}}{\lambda_{i}+\varepsilon}$ s.t. we recover the marginal constraint for $\tau_{i}=1$, when $\lambda_{i}\rightarrow+\infty$. We write $\tau=(\tau_{1},\tau_{2})$.
@@
Flow Matching (FM). Given a prior $\rho_{0}\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$ and a time-dependent vector field $(v_{t})_{t\in[0,1]}$, one can define a probability path $p_{t}$ starting from $\rho_{0}$ using the flow $\phi_{t}$ solving
@@
by setting $p_{t}=\phi_{t}\sharp\rho_{0}$. We then say that $v_{t}$ generates the path $p_{t}$. Continuous Normalizing Flows [9] (CNFs) model $v_{t,\theta}$ with a neural network , which is trained to match a terminal condition $p_{1}=\rho_{1}\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$. FM [45] is a simulation-free technique to train CNFs by constructing individual paths between samples, and minimizing
@@
where $[Z_{0},Z_{1}]_{t}:=(1-t)Z_{0}+tZ_{1}$ with $t\sim\mathcal{U}([0,1])$. If this loss is 0, the flow maps $\rho_{0}$ to $\rho_{1}$, i.e., $\phi_{1}\sharp\rho_{0}=\rho_{1}$. This property is often referred to as FM preserving the marginal distribution.